OpenStudy (anonymous):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 9 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (sleepyjess):

do you know how to find the volume of the cylinder and cone?

OpenStudy (anonymous):

maybe... @sleepyjess

OpenStudy (sleepyjess):

do you know the formulas?

OpenStudy (anonymous):

\[v = 3.14 r^2 h\]

OpenStudy (tkhunny):

Okay, now look up the cone.

OpenStudy (anonymous):

it says its a 3-d shape with a flat bottom...

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

ok you have the formula for the volume of the cylinder, can you find the volume?

OpenStudy (anonymous):

thats what i need help with @sleepyjess

OpenStudy (sleepyjess):

\(v=3.14*r^2*h\) r=diameter/2

OpenStudy (anonymous):

so about 452

OpenStudy (sleepyjess):

yes

OpenStudy (sleepyjess):

for a cone \(V=3.14*r^2*\dfrac{h}{3}\)

OpenStudy (sleepyjess):

just do the same thing with the cone as you did for the cylinder

OpenStudy (anonymous):

so 1206

OpenStudy (sleepyjess):

can you show me what you did to get to that?

OpenStudy (anonymous):

i plugged the 8 and 18 into that formula you gave me

OpenStudy (sleepyjess):

remember r=diameter/2

OpenStudy (anonymous):

radius is half of diameter?

OpenStudy (anonymous):

So about 302

OpenStudy (anonymous):

thanks.. gave you the medal you deserved

OpenStudy (sleepyjess):

it would actually be 301 since the real value is 301.44.

OpenStudy (anonymous):

tk can you help me again?